This is a weighted average problem. Hope I chose the right section....
boat number one leaves the pier at 9am at 8 knots. 1/2 hour later boat number two leaves the same pier in the same direction at 10 knots. What time will boat two overtake boat number one? I have examples in my book but they are going in different directions and it doesn't work with the same type of formula for me. Thank you!!!!

Answer by checkley71(8403)

THE FORMULA IS 8T=10(T-.5) WHERE T IS THE TIME WHEN THE BOATS WILL MEET.
8T=10T-5 OR 5=10T-8T OR 5=2T THEREFORE T=2.5 HOURS.
8 X 2.5=20 & 10 X 2=20.